OpenStudy (anonymous):

Solve. 2x2 = 8 x = 2 x = 2 and x = –2 x = –2 and x = 4 x = 4

OpenStudy (mertsj):

Divide both sides by 2 first.

OpenStudy (anonymous):

which is 2 and 4?

OpenStudy (mertsj):

\[\frac{2x^2}{2}=\frac{8}{2}\]

OpenStudy (mertsj):

\[x^2=4\]

OpenStudy (mertsj):

\[x=\pm 2\]